An arrow is fired into the air with an initial velocity of 64 feet per second. The height in feet of the arrow after t seconds after it was shot into the air is given by the function h(x)=-16t(squared) +64t. Find the maximum height of the arrow.
Answer by lwsshak3(11628)

An arrow is fired into the air with an initial velocity of 64 feet per second. The height in feet of the arrow after t seconds after it was shot into the air is given by the function h(x)=-16t(squared) +64t. Find the maximum height of the arrow.
..
h(x)=-16t(squared) +64t.
y=-16t^2+64t
Change equation into standard form for parabola: y=A(x-h)^2+k, with (h,k) being the (x,y) coordinates of the vertex.
completing the square
y=-16(t^2-4t+4)+64
y=-16(t-2)^2+64
vertex(2,64)
ans:
The arrow will reach its maximum height of 64 feet 2 seconds after it was shot into the air.